# MRAG: A Modular Retrieval Framework for Time-Sensitive Question Answering

### v1

@@ 5.1 Experimental Setup
For retrieval, we include BM25, Contriever, and a hybrid method (Jedidi et al., 2024). Reranking methods include ELECTRA (Clark et al., 2020), MiniLM (Wang et al., 2020), Jina(Jina, 2024), BGE (Xiao et al., 2023), NV-Embed (Lee et al., 2024), and Gemma (Gemma et al., 2024). We follow state-of-the-art answer generation approaches based on prompting LLMs (§2.2). We evaluate four approaches, Direct Prompt that adopts question-answer pairs as few-shot examples; Direct CoT that adds rationals into prompts (Wei et al., 2022); RAG-Concat, where passages are concatenated into a LLM; and Self-RAG, which processes each passage independently and selects the best answer (Asai et al., 2024).
@@
Due to limited budget, we evaluate GPT-4o mini (OpenAI et al., 2024) with direct prompting, and three open-source LLMs: TIMO, a LLaMA2-13B model fine-tuned for temporal reasoning (Su et al., 2024b), and two general-purpose models, Llama3.1-8B-Instruct and Llama3.1-70B-Instruct Dubey et al. (2024). We use 10 examples in prompts.
@@
Table 2: The answer recall (AR@k) and gold evidence recall (ER@k) of each retrieval method on perturbed temporal queries in TimeQA and SituatedQA subsets of TempRAGEval. 1st means the first-stage retrieving method; 2nd means the second-stage reranking method; # QFS means the number of top passages to be summarized. Bold numbers indicate the best performance. We include complete results in Appendix E.
@@ 5.2 Main Results
Table 3: End-to-end QA comparison for various generation strategies and LLMs on TempRAGEval. Bold numbers indicate the best performance the each backbone LLM. The second best is underlined.
@@
As shown in Table 3, MRAG outperforms baseline RAG approaches in QA accuracy, for instance 49.2% EM (MRAG) over 44.0% EM (RAG) in TempRAGEval-TimeQA for Llama3.1-8B. This highlights the role of enhanced retrieval in improving accuracy of the time-sensitive QA task. Incorporating a self-reflection strategy improves performance for Llama3.1 models but not for TIMO. We observe that TIMO makes more errors in self-reflection, likely due to the limited reasoning capacity of its backbone model, Llama 2.
@@ 6 Analysis
In this part, we provide detailed analysis on how each component of MRAG contributes to the overall improvement.
@@ 6.1 Ablation Study
The impact of the numbers of passages for summarization.
Our LLM based summarization removes irrelevant temporal information but may also introduce hallucinations. We conduct a parametric study to evaluate the impact of the number of passages used for summarization. As shown at the bottom of Table 2, summarizing top-five passages achieves the best balance.
The impact of the number of passages used for answer generation in MRAG.
Our experiments show that the optimal number of passages depends on the size of the LLM. TIMO can handle a maximum of 3 passages, while the optimal number for LLaMA 3.1-8B is five, and for LLaMA 3.1-70B, it is twenty. Larger LLMs, with more parameters, process longer contexts better. Full results are presented in LABEL:sec:_concat.
@@ 6.2 Human Evaluation
One limitation of the retrieval metrics is that AR@k overestimates performance, as a passage might incidentally contain an answer without directly supporting it. Conversely, ER@k acts as a conservative lower bound, potentially overlooking other relevant but unannotated passages. To address this, we conduct a human evaluation on a random subset of 200 examples to assess actual retrieval performance. We manually review the top 20 passages retrieved by MRAG and Gemma, annotating the first relevant passage. The results confirm that MRAG consistently outperforms Gemma in retrieval accuracy. We present full results in LABEL:sec:_human.
@@ 6.3 Case Study
Figure 4: A case study for top-1 passages retrieved by Gemma and MRAG for a question from TempRAGEval-SituatedQA.
We conduct case studies to qualitatively assess the advantages of MRAG over the Gemma baseline. According to Figure 4, the top-1 passage by Gemma matches the query date “1988”, but discusses a father-son record set in 2007. In contrast, MRAG result is about a teammate combination record in the same season, despite the date “1961” differing from the query date “1988”. We include more cases in LABEL:sec:_cases.
@@ 7 Other Related Works
Recent research has explored LLM embeddings for retrieval. Some studies focus on distilling or fine-tuning LLM embeddings for reranking tasks, such as Gemma (Gemma et al., 2024) and MiniCPM (Hu et al., 2024). Others aim to develop generalist embedding models capable of performing a wide range of tasks including retrieval and reranking, e.g., gte-Qwen (Yang et al., 2024) and NV-Embed (Lee et al., 2024). These LLM-based methods have demonstrated unprecedented performance in benchmarks such as MTEB (Muennighoff et al., 2023) over Contriver-based dense retrievers, and also on our TempRAGEval benchmark.
@@
Existing retrieval benchmarks primarily target keyword-based or semantic-based retrieval. Su et al. (2024a) introduces BRIGHT, a new retrieval benchmark emphasizing intensive reasoning. The benchmark challenges retrievers to perform reasoning across diverse domains, such as analogies and algorithms, to identify relevant documents. We focus on temporal reasoning-intensive retrieval, sharing a similar vision for future advanced retrieval systems.
@@ 8 Conclusion
This study focuses on time-sensitive QA, a task that challenges LLM based QA systems. We first present TempRAGEval, a diagnostic benchmark featuring natural questions, evidence annotations, and temporal perturbations. We further propose a training-free MRAG framework, which disentangles relevance-based retrieval from temporal reasoning and introduces a symbolic temporal scoring mechanism. While existing systems struggle on TempRAGEval due to limited temporal reasoning capacities in retrieval, MRAG shows significant improvements. We hope this work advances future research on reasoning-intensive retrieval.
@@
There are still some limitations in our work: (1) Our proposed benchmark is designed to evaluate time-sensitive questions with explicit temporal constraints. However, addressing questions with implicit temporal constraints presents a more complex challenge for retrieval systems. Our symbolic approach is not directly applicable to such questions and requires more advanced reasoning mechanisms to bridge this gap. (2) Our benchmark primarily targets time-sensitive questions with a single answer. Future work can extend to predict a list of answers with different temporal constraints. (3) Despite our approach improves temporal reasoning in retrieval, there is still head room. Fine-tuning dense retrieval models on temporal reasoning-intensive retrieval datasets may offer a promising solution for future work.
@@ Appendix A Controlled Experiments
In summary, existing retrieval methods demonstrate limited temporal reasoning capabilities. The LLM embedding-based method shows better performance than others.
@@
Appendix B Implementation Details
We conduct empirical evaluations for MRAG and SOTA retrieving-and-reranking systems on TempRAGEval. MRAG consists of functional modules, which can be based on algorithms, models, or prompting methods. In implementation, algorithm based modules include question normalization, keyword ranking, time extraction, and semantic-temporal hybrid ranking. Model based modules are retrieving, semantic ranking, sentence tokenization. To ensure fair comparison, we use the same retriever model, i.e., Contriever (Izacard et al., 2021), as the first stage method for MRAG and two-stage systems. We use Gemma embeddings (Li et al., 2023) as the main tool to measure semantic similarity for passages and sentences in MRAG. NLTK package is used for sentence tokenization (Bird and Loper, 2004). LLM prompting based modules are keyword extraction, query-focused summarization. As shown in Appendix E, we have tested Llama3.1-8B-Instruct and Llama3.1-70B-Instruct models (Dubey et al., 2024) for LLM prompting based modules. Detailed prompts are listed in LABEL:sec:_prompts. The evaluation metrics are computed based on retrieved passages not sentences.
Appendix C Retrieval Performance Degradation from Perturbations
We compare the retrieval performance between original queries and perturbed queries using the same baseline retrieval system (i.e., Contriever retriever and Gemma reranker). For both the TimeQA and SituatedQA subsets, the perturbed questions significantly increase the difficulty of retrieving relevant documents, particularly when evaluating the top-1 and top-5 ranked documents. This suggests that the introduction of perturbations introduces greater complexity. The existing retrieval method has limited temporal reasoning capabilities and is not robust to such variations.
@@ Appendix C Retrieval Performance Degradation from Perturbations
Appendix D Spline Functions for Temporal Scoring
In MRAG, the temporal relevance is determined symbolically using temporal constraints in query, extracted dates from sentences, and pre-defined scoring functions. The temporal constraints are categorized by implicit conditions, i.e., “first” and “last”, and temporal relations, i.e., “before”, “after”, and “between”. For example, the question “Who won the latest game as of 1981?” will be classified as the “last” type and the “before” type. We pre-define spline function templates for each type of temporal constraints, as shown in Figure 7. Each spline function computes a temporal score for a given input date extracted from the sentence. For sentences with multiple dates, the highest temporal score is selected.
@@ Appendix D Spline Functions for Temporal Scoring
Appendix E Complete Retrieval Evaluation Results

### v2

@@ 5.1 Experimental Setup
For retrieval, we include BM25, Contriever, and a hybrid method (Jedidi et al., 2024). Reranking methods include ELECTRA (Clark et al., 2020), MiniLM (Wang et al., 2020), Jina(Jina, 2024), BGE (Xiao et al., 2023), NV-Embed (Lee et al., 2024a), and Gemma (Gemma et al., 2024). We follow state-of-the-art answer generation approaches based on prompting LLMs (§2.2). We evaluate four approaches, Direct Prompt that adopts question-answer pairs as few-shot examples; Direct CoT that adds rationals into prompts (Wei et al., 2022); RAG-Concat, where passages are concatenated into a LLM; and Self-RAG, which processes each passage independently and selects the best answer (Asai et al., 2024).
@@
Due to limited budget, we evaluate GPT-4o mini (OpenAI et al., 2024) with direct prompting, and three open-source LLMs: TIMO, a LLaMA2-13B model fine-tuned for temporal reasoning (Su et al., 2024b), and two general-purpose models, Llama3.1-8B-Instruct and Llama3.1-70B-Instruct Dubey et al. (2024). We use 10 examples in prompts. To eliminate the impact of input length constraints across models, we conduct parametric studies as described in LABEL:sec:_concat and report each LLM’s performance with its optimal number of input passages in Table 3.
@@
Table 2: The answer recall (AR@k) and gold evidence recall (ER@k) of each retrieval method on perturbed temporal queries in TimeQA and SituatedQA subsets of TempRAGEval. 1st means the first-stage retrieving method; 2nd means the second-stage reranking method; # QFS means the number of top passages to be summarized. Bold numbers indicate the best performance. We include complete results in Appendix I.
@@ 5.2 Main Results
Table 3: End-to-end QA performance comparison for various generation strategies and LLMs on TempRAGEval. Bold numbers indicate the best performance the each backbone LLM. The second best is underlined. TIMO has a limited input length with up to three passages. We report the best number of passages for Llama models, and provide ablation on different numbers in LABEL:sec:_concat.
@@
As shown in Table 3, MRAG outperforms baseline RAG approaches in QA accuracy, for instance 49.2% EM (MRAG) over 44.0% EM (RAG) in TempRAGEval-TimeQA for Llama3.1-8B. Incorporating a self-reflection strategy improves performance for Llama3.1 models but not for TIMO, likely due to the limited reasoning capacity of its backbone model, Llama 2.
@@ 6 Analysis
This section presents a detailed analysis of the results and the contribution of each MRAG module.
@@ 6.1 Ablation Study
The impact of the number of passages for summarization.
Our LLM based summarization removes irrelevant temporal information but may also introduce hallucinations (details in Appendix H). The parametric study at the bottom of Table 2 shows that summarizing top-five passages achieves the best balance. Additionally, we compare the RAG setup with the long-document QA setup bypassing retrievers in LABEL:longdoc. RAG achieves a better accuracy, as retrievers exclude irrelevant passages and reduce noise.
The impact of the number of passages for answer generation.
Our experiments show that the optimal number of passages depends on the LLMs. TIMO can handle a maximum of 3 passages, while the optimal number for LLaMA 3.1-8B is five, and for LLaMA 3.1-70B, it is twenty. Full results are presented in LABEL:sec:_concat.
The impact of key retrieval steps in MRAG.
In the retrieval and summarization module, with retrieved passages, MRAG conducts two key steps: (1) Passage Keyword Ranking reduces the passages from 1,000 to 100 based on the keyword presence; (2) Passage Semantic Ranking reorders the top-100 passages and splits them into $\sim$500 sentences, including chunk summaries. Similarly, in the semantic-temporal hybrid ranking module, another two key steps are conducted: (1) Sentence Keyword Ranking further narrows the scope to 200 sentences; (2) Hybrid Ranking ranks these sentences by semantic-temporal hybrid scoring. As shown in LABEL:ablation, keyword-based ranking steps effectively reduce ranking candidates without significant performance loss, while semantic and hybrid ranking steps markedly enhance retrieval performance. An efficiency-accuracy trade-off can be made by adjusting the number of targeted passages or sentences at each step.
Computational Overhead.
As MRAG involves retrieval, summarization, and re-ranking, it incurs approximately twice the computational overhead of standard RAG pipelines, which is manageable. We provide a detailed analysis in LABEL:overhead.
@@ 6.2 Human Evaluation
One limitation of the retrieval metrics is that AR@k overestimates performance, as a passage might incidentally contain an answer without directly supporting it. Conversely, ER@k acts as a conservative lower bound, potentially overlooking other relevant but unannotated passages. To address this, we conduct a human evaluation on a random subset of 200 examples to assess actual retrieval performance. The results validate the advantages of MRAG over Gemma in retrieval accuracy with full results presented in LABEL:sec:_human.
@@ 6.3 Case Study
Figure 4: A case study for top-1 passage retrieved by Gemma and MRAG from TempRAGEval.
We analyze retrieval errors qualitatively to highlight the advantages of MRAG over the Gemma baseline. As shown in Figure 4, the top-1 passage by Gemma matches the query date “1988” but discusses a father-son record set in 2007. In contrast, MRAG retrieves a passage about a teammate combination record from the same season, despite the differing date (“1961” vs. “1988”). Additional cases and answer generation error cases are provided in LABEL:sec:_cases_retrieval and LABEL:sec:_cases_qa.
@@ 7 Other Related Works
Recent research has explored LLM embeddings for retrieval. Some studies focus on distilling or fine-tuning LLM embeddings for reranking tasks, such as Gemma (Gemma et al., 2024) and MiniCPM (Hu et al., 2024). Others aim to develop generalist embedding models capable of performing a wide range of tasks including retrieval and reranking, e.g., gte-Qwen (Yang et al., 2024) and NV-Embed (Lee et al., 2024a). These LLM-based methods have demonstrated unprecedented performance in benchmarks such as MTEB (Muennighoff et al., 2023) and our TempRAGEval.
@@
Existing retrieval benchmarks primarily target keyword-based or semantic-based retrieval. Su et al. (2024a) introduces BRIGHT, a new retrieval task emphasizing intensive reasoning. We focus on temporal reasoning, one aspect of a broader class of reasoning-intensive retrieval. MRAG is expected to generalize to other forms of symbolic reasoning, such as numeric ranges and geospatial constraints. It mitigates direct numeric matching in retrieval and enhances reasoning capabilities.
@@ 8 Conclusion
This study focuses on time-sensitive QA, a task that challenges LLM based QA systems. We first present TempRAGEval, a diagnostic benchmark featuring natural questions, evidence annotations, and temporal complexity. We further propose a training-free MRAG framework, which disentangles relevance-based retrieval from temporal reasoning and introduces a symbolic temporal scoring mechanism. While existing systems struggle on TempRAGEval due to limited temporal reasoning capacities in retrieval, MRAG shows significant improvements. We hope this work advances future research on reasoning-intensive retrieval.
@@
There are still some limitations in our work: (1) Our proposed benchmark is designed to evaluate time-sensitive questions with explicit temporal constraints. However, addressing questions with implicit temporal constraints presents a more complex challenge for retrieval systems. We could extend to implicit temporal reasoning by associating explicit information to the implicit one using LLM common sense and background knowledge like (Chen et al., 2022). (2) Our dataset does not include time-sensitive questions that fall outside the LLM knowledge cutoff. We could extend our dataset with questions from RealTime QA (Kasai et al., 2024) and AntiLeak-Bench (Wu et al., 2024b). (3) Our main objective is to improve temporal reasoning in retrieval, which has not been tackled by previous works. More complex scenarios like multi-hop and recursive reasoning require further research efforts. (4) The proposed framework introduces computational overhead for improved performance as detailed in LABEL:overhead. (5) We analyze knowledge conflicts between LLMs and passages in LABEL:sec:_conflict and leave conflicts among passages for future work.
Ethics Statement
TempRAGEval were constructed upon the test set of TimeQA (Chen et al., 2021) and SituatedQA (Zhang and Choi, 2021) datasets, which are publicly available under the licenses BSD-3-Clause license [^8] https://opensource.org/licenses/BSD-3-Clause and Apache-2.0 license [^9] https://www.apache.org/licenses/LICENSE-2.0. These licenses all permit us to compose, modify, publish, and distribute additional annotations upon the original dataset. All the experiments in this paper were conducted on 4 NVIDIA L40S 46G GPUs. We hired 3 graduate students in STEM majors as annotators. We recommended that annotators spend at most 2 hours per day for annotation in order to reduce pressure and maintain a comfortable pace. The whole annotation work lasted about 5 days.
@@ Appendix A Controlled Experiments
In summary, existing retrieval methods demonstrate limited temporal reasoning capabilities. The LLM embedding-based method shows better performance than others. Our proposed MRAG framework is retriever-agnostic, which aims to improve temporal reasoning capabilities for any type of retrieval models.
@@
Appendix B Dataset Selection Criteria
Our benchmark focuses on time-sensitive question answering, which is knowledge-intensive. Therefore, datasets designed for only temporal reasoning (e.g., “What is the time 5 year and 5 month after Oct, 1444”) are not considered, such as TempReason (Tan et al., 2023) and DurationQA (Virgo et al., 2022). Aggregated benchmarks (e.g., TimeBench (Chu et al., 2024) and TRAM (Wang and Zhao, 2024)) focus on evaluating diverse temporal reasoning capabilities, which have a broader scope than our focus. MenatQA (Wei et al., 2023) is built by adding counterfactual and order factors to TimeQA (Chen et al., 2021) questions, which is similar to our approach. Other knowledge-intensive temporal QA datasets can serve as alternative sample sources including StreamingQA Liška et al. (2022), TempLAMA (Dhingra et al., 2022), and concurrent dataset ComplexTQA (Gruber et al., 2024). We select SituatedQA (Zhang and Choi, 2021) for its human-written questions, distinguishing it from other temporal QA datasets that are typically synthetic. Additionally, we opt for TimeQA (Chen et al., 2021) due to its hard split, which already includes complex temporal questions. Notably, both SituatedQA and TimeQA can be grounded in the Wikipedia corpus.
Appendix C Annotation Guidelines
C.1 Annotating Perturbations
Given a question-answer pair sourced from TimeQA or SituatedQA (e.g., Q: “Arnolfini Portrait was owned by whom between Jul 1842 and Nov 1842?” A: “National Gallery”), annotators should ground the pair to facts in Wikipedia (e.g., “The Arnolfini Wedding by Jan van Eyck, has been part of the National Gallery’s collection in London since 1842.”). Then they identify the key timestamps or durations of Wikipedia facts (e.g., 1842). To create temporal perturbations, annotators are asked to come up with combinations of implicit conditions, temporal relations, and alternative dates to form complex temporal constraints. The implicit condition can be “None” or selected from a list of 4 types: “first”, “earliest”, “last”, and “latest”. The temporal relation should be selected from a list of 11 types: “as of”, “from to”, “until”, “before”, “after”, “around”, “between”, “by”, “in”, “on”, and “since”. Finally, annotators rewrite questions naturally (e.g., “Who is the last one owned Arnolfini Portrait after 1700?”) by introducing perturbed temporal constraints (e.g., “last … after 1700”) and ensure that the answers (e.g., “National Gallery”) remain unchanged. After different annotators create perturbed question-answer pairs, they exchange these pairs with each other to validate the correctness of the answers. Only the perturbed samples validated by two annotators are kept.
C.2 Annotating Gold Evidences
For gold evidence annotations, annotators are assigned different perturbed question-answer pairs. For each pair, 20 context passages are provided to annotators, which are retrieved by the leading retriever Contriever (Izacard et al., 2021), and the best reranker Gemma (Li et al., 2023). Annotators are asked to identify up to two gold evidence passages from these passages. A passage is regarded as relevant and annotated as gold evidence if annotators can obtain the correct answer from this passage. If there is no relevant passage among these 20 retrieved ones, annotators should search Wikipedia pages related to the query entities to locate the gold evidence passages manually. Lastly, annotators exchange samples to validate gold evidence annotated by others. Only the gold evidence annotations validated by two annotators are kept.
Appendix D Sample Statistics
We gather a similar size of examples as previous temporal QA benchmarks (e.g., 3K for TimeQA (Chen et al., 2021) and 2K for MenatQA (Wei et al., 2023)), which is enough for an evaluation set (see examples in LABEL:examples). As we manually annotate gold evidence passages in Wikipedia, it is time-consuming to scale up like other synthetic datasets (Dhingra et al., 2022; Gruber et al., 2024). To understand the difference between two subsets, we summarize the statistics in Appendix D. The average length of questions is measured by the GPT-2 tokenizer (Radford et al., 2019). We assess the popularity of key entities in questions using the average monthly page view counts of the corresponding Wikipedia page in 2024 (Pageviews, 2024). As we can see, the main difference lies in the question entity popularity. TempRAGEval-TimeQA questions typically inquire about lesser-known individuals and are generally straightforward and clear. In contrast, TempRAGEval-SituatedQA questions commonly ask about a sports team and championship, which are more general and sometimes ambiguous. This difference may explain varying retrieval and QA performance across the two subsets.
TimeQA SituatedQA
# original questions 123 120
# perturbed questions 377 380
# total questions 500 500
Temporal complexity hard hard
Avg. question length 15.2 15.6
Avg. entity popularity 7,456 57,521
Table 4: Sample statistics for TempRAGEval-TimeQA and TempRAGEval-SituatedQA.
@@ Appendix D Sample Statistics
Appendix E Retrieval Performance Degradation Due to Perturbations
As shown in Figure 6, we compare the retrieval performance between original queries and perturbed queries using the same baseline retrieval system (i.e., Contriever retriever and Gemma reranker). For both the TimeQA and SituatedQA subsets, the perturbed questions significantly increase the difficulty of retrieving relevant documents, particularly when evaluating the top-1 and top-5 ranked documents. This suggests that the introduction of perturbations introduces greater complexity. The existing retrieval method has limited temporal reasoning capabilities and is not robust to such variations.
Appendix F Evaluation Experiment Implementation Details
We conduct empirical evaluations for MRAG and SOTA retrieving-and-reranking systems on TempRAGEval. In baselines, due to limited computing resources, we use LLM-based embedding models as a reranking model, such as Gemma (Gemma et al., 2024) and NV-Embed (Lee et al., 2024a). MRAG consists of functional modules, which can be based on algorithms, models, or prompting methods. In implementation, algorithm based modules include question normalization, keyword ranking, time extraction, and semantic-temporal hybrid ranking. Model based modules are retrieving, semantic ranking, sentence tokenization. To ensure fair comparison, we use the same retriever model, i.e., Contriever (Izacard et al., 2021), as the first stage method for MRAG and two-stage systems. We use Gemma embeddings (Li et al., 2023) as the main tool to measure semantic similarity for passages and sentences in MRAG. NLTK package is used for sentence tokenization (Bird and Loper, 2004). LLM prompting based modules are keyword extraction, query-focused summarization. As shown in Appendix I, we have tested Llama3.1-8B-Instruct and Llama3.1-70B-Instruct models (Dubey et al., 2024) for LLM prompting based modules. Detailed prompts are listed in LABEL:sec:_prompts. The evaluation metrics are computed based on retrieved passages not sentences.
@@ Appendix F Evaluation Experiment Implementation Details
Appendix G Implementations for Semantic-Temporal Hybrid Scoring
The Retrieval and Summarization module in MRAG splits and summarizes top relevant passages into independent sentences for the downstream fine-grained reranking, which is inspired by Yang (2023). The Semantic-Temporal Hybrid Scoring module is designed to assess the semantic relevance and temporal relevance between the question and the evidence sentence. To quantify the semantic relevance, we apply an embedding model (e.g., Gemma) to the question main content and the sentence.
For temporal relevance, we employ a symbolic scoring approach, wherein the module automatically generates scoring functions for each question and computes temporal scores for individual sentences.
To generate scoring functions, we classify question temporal constraints into six categories and define a template for each. A scoring function is instantiated using the corresponding template and extracted timestamp(s) from the question. The six constraint types include: “first - before”, “first - after”, “first - between”, “last - before”, “last - after”, and “last - between”. Here, “last” denotes that the question seeks the most recent event, while “before” indicates that the event must precede a specified date. For instance, the question “Who won the latest game as of 1981?” corresponds to “last - before - 1981”, as illustrated in the top-left subplot of Figure 7.
Afterwards, the module extracts the timestamp(s) from the evidence sentence. It computes the temporal score based on the extracted timestamp and the corresponding scoring function. If multiple timestamps are present, the highest temporal score is selected. For example, as depicted in Figure 7, for the constraint “last - before - 1981”, an evidence sentence mentioning “1970” would receive a temporal score around 0.9.
The final score of each evidence sentence for each question is obtained by multiplying the temporal score and the semantic score. Finally, we select the passages that contains the highest-scoring sentences. The passages are fed into the later answer generation stage rather than the sentences. The passages provide better background information, which leads to higher generation quality for reader systems.
Appendix H LLM-based Summarization Case Study
LLM-based query-focused summarization enhances retrieval performance by distilling key information from passages while filtering out irrelevant context, as demonstrated in Table 5. In the second success case, the summarization effectively converts structured data into natural language, benefiting retrievers that are primarily trained on free-text retrieval tasks. However, LLM-generated summaries may introduce hallucinations and errors, though their occurrence is infrequent. As shown in Table 5, erroneous summaries can mislead the retriever with non-factual events or incorrect dates, resulting in irrelevant passages ranking higher. To balance retrieval improvements with the risk of errors, we summarize only the top-$k$ passages per query, which also reduces computational overhead. Furthermore, to prevent error propagation, we provide the reader model with original passages rather than their summaries.
Success Cases
Question Who won the latest America’s Next Top Model by May 8, 2021?
Answer Kyla Coleman
Passage America’s Next Top Model (season 24) | The twenty-fourth cycle of America’s Next Top Model premiered on January 9, 2018 … The winner of the competition was 20 year-old Kyla Coleman from Lacey, Washington with Jeana Turner placing as the runner up.
Summarization Kyla Coleman, a 20-year-old from Lacey, Washington, won the competition in 2018.
Question When was the last time Kentucky won NCAA in basketball after 2010?
Answer 2012
Passage Kentucky Wildcats | Men (8) ; Basketball (8): 1948, 1949, 1951, 1958, 1978, 1996, 1998, 2012 ; Women (2) … List of NCAA schools with the most NCAA Division … Kentucky has won 13 NCAA team national championships.
Summarization The Kentucky Wildcats won the NCAA basketball championship in 1948, 1949, 1951, 1958, 1978, 1996, 1998, and 2012.
Failure Cases
Question When was the last time the Ducks won the Stanley Cup as of 2010?
Answer 2007
Passage Anaheim Ducks | … Despite the arenas being six hours away from each other, the teams have developed a strong rivalry, primarily from the 2009 and 2018 Stanley Cup playoffs. The Ducks won the series in 2009, but the Sharks came back in 2018.
Summarization The Anaheim Ducks won the Stanley Cup in 2009.
True fact The Anaheim Ducks won the Stanley Cup in 2007. That was their first and only championship so far.
Question How many times has South Korea held the Winter Olympics as of 2018?
Answer 1 | one
Passage 2018 Winter Olympics | The 2018 Winter Olympics … This marked the second time that South Korea had hosted the Olympic Games (having previously hosted the 1988 Summer Olympics in Seoul) …
Summarization South Korea held the Winter Olympics in 2018 and previously in 1988.
True fact South Korea held the Winter Olympics in 2018 and Summer Olympics in 1988.
Table 5: Success and error cases of LLM-based query-focused summarization using Llama3.1-8B-Instruct.
Appendix I Complete Retrieval Evaluation Results
We evaluate MRAG on TempRAGEval with baseline retrieval methods, including ELECTRA101010cross-encoder/ms-marco-electra-base, MiniLM 111111cross-encoder/ms-marco-MiniLM-L-12-v2, Jina121212jinaai/jina-reranker-v2-base-multilingual, BGE [^13] BAAI/bge-reranker-large, NV-Embed141414nvidia/NV-Embed-v1, and Gemma  [^15] BAAI/bge-reranker-v2-gemma. Complete rsults are presented in LABEL:experiments_timeqa and LABEL:experiments_situatedqa.